OpenStudy (anonymous):

4 sq rt (z-3)=3 sq rt (z+4)

OpenStudy (kainui):

I'd probably square both sides of this equation. Remember, for an equation to be true both things one either side have to be exactly the same. So you must do exactly the same thing to both sides every time you do something to keep the equality true.

OpenStudy (anonymous):

z=24?

OpenStudy (kainui):

Almost, it's a multiple of that. Maybe you messed up a 2 somewhere? Check your work or write it out and I'll show you where you went wrong so you don't make that mistake again. =)

OpenStudy (anonymous):

after I square both sides the equation I have is 4z-12=3z+12 z-12=12 z=24

OpenStudy (kainui):

\[4 \sqrt {z-3}=3 \sqrt {z+4}\]\[(4 \sqrt {z-3})^2=(3 \sqrt {z+4})^2\]\[4^2 \sqrt {z-3}^2=3^2 \sqrt {z+4}^2\]\[16( z-3)=9(z+4)\] Can you follow my reasoning here? I think you just forgot to square those coefficients outside the square root signs as well, that's all.

OpenStudy (anonymous):

so its 12?